OpenStudy (anonymous):

HELP!!!!!!! Grant can row 6 miles in 36 minutes. Arianna can row 3,520 yards in 20 minutes. In miles per hour, how much faster does grant row than Arianna ? (note: 1 mile=1,760 yard)

OpenStudy (anonymous):

12times faster

OpenStudy (phi):

can you change 3520 yards to miles ? can you change 20 minutes to hours ?

OpenStudy (anonymous):

yea.

OpenStudy (anonymous):

how are you supposed to change 20min in to hours

OpenStudy (anonymous):

lol

OpenStudy (phi):

to change yards to miles, think like this: we have yards, we want miles so we want to do \[ \text{yards} \cdot \frac{\text{miles}}{\text{yards}} \] so that the yards cancel, and you get \[ \cancel{\text{yards}} \cdot \frac{\text{miles}}{\cancel{\text{yards}} } = \text{miles}\]

OpenStudy (anonymous):



OpenStudy (anonymous):